### v2

@@ 4. Volumetric Primitives
Inspired by recent works using Gaussian primitives for reconstructing radiance fields (Kerbl et al., 2023), and following a similar medium definition as Knoll et al. (2021), we model media using sets of primitives. Each primitive $\mathcal{P}_{i}$ represents a statistical aggregate of matter with identical emission, cross section and phase function, and with density $\rho_{i}(\mathbf{x})$ defined by a three-dimensional un-normalized kernel $K_{i}(\mathbf{x})$. Given these primitives, we can model the distribution of matter, and therefore the extinction probability as a mixture of these volumetric primitives following
@@
with $N$ the number of primitives affecting $\mathbf{x}$, and $\sigma_{i}$ its cross-section. For kernels with infinite support, $N$ is the total number of primitives $N_{\text{total}}=N$. In our case, we assume that all kernels have limited support, so typically $N<N_{\text{total}}$. We compute the single-scattering albedo and phase function at $\mathbf{x}$ analogously, with the key difference of requiring normalization. Details can be found in the supplementary material.
@@ 4.1. The RTE with volumetric primitives
Defining matter as a mixture of volumetric primitives with finite support allow us to reinterpret Equation (3) as the summation of the contribution of all the primitives. For that, as shown in Figure 2 (bottom), we split the ray in segments based on the primitives’ entry and exit points; then, we integrate radiance along segments in a front-to-back pass, keeping track of the multiplicative transmittance which is computed analytically. This approach is similar to how production renderers deal with complex overlapping media (Fong et al., 2017, Ch.6); the key difference is that our approach approximates the media with primitives, which allows having closed-form transmittance and sampling expressions.
@@ 4.2. Solving Equation 7 for distance sampling
While Equation (7) is just a summation over segments that can be computed analytically (we show examples later in Section 7), in the most general case $L_{o,i}(\mathbf{x}_{t},\omega)$ hides the inscattering integral, which requires numerical evaluation using Monte Carlo sampling. Computing a Monte Carlo estimate would quickly become impractical; thus we need to sample a single segment per ray, which requires essentially sampling the distance $t$ along the ray, ideally with probability distribution function (PDF) $p(t)=\mu_{t}(\mathbf{x}_{t})\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})$.
While our primitive-based media allows the use of delta tracking (or any other Monte Carlo-based) estimators for distance sampling, we found that we can leverage the recursive formulation of transmittance (9) and pose the problem as an iterative search problem, where we uniformly sample the transmittance with a random variable $\xi\in(0,1)$, and search for the distance $t$ so that $\xi(t)=\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})$. This approach is very similar to regular tracking (Sutton et al., 1999), which we extend to handle overlapping kernels efficiently. In particular, we first search for a segment $\mathcal{S}_{k}$ so that $(1-\xi)\in[\textrm{T}_{k-1},\textrm{T}_{k}]$. Then, we invert the transmittance inside the segment, solving for $t\in[\widehat{t}_{k,0},\widehat{t}_{k,1}]$ the following equation:
@@
Depending on the type of kernel $K_{i}(\mathbf{x})$ being used, and the amount of overlaps in segment $\mathcal{S}_{k}$ this equation may have a closed-form analytical solution. However, for cases where Equation (12) does not have a simple analytical solution (e.g. many overlapping kernels at the same segment), we rely on numerical root-finding. In particular, we use two different methods depending on the kernel: the Newton-Raphson and the bisection solvers. The former is our primary choice, since it is efficient and converges with good precision with a few iterations (1-3), given that our problem is well-conditioned. However, for certain long-tailed kernels (e.g., Gaussian) it occasionally suffers from numerical instability. In these cases, we rely in the slightly slower but more stable bisection method. Details for both methods can be found in the Supplemental. Additionally, to maximize speed, for complex assets fitted with a high number of small kernels, we can avoid inversion altogether and simply sample the segment uniformly. This introduces a small amount of bias in exchange for a significant performance boost, which we analyze in Section 7.1.
Discussion
We have derived a general framework for radiative transport in primitive-based media, defined using arbitrary kernel functions. Any kernel could be used within our framework as long as 1) they have limited support or their decay is such that they can be bounded or clipped efficiently and 2) closed-form solutions to their line integrals exist or can be numerically computed efficiently.
@@ 5. Kernels
We implement our framework using two different kernels: the Gaussian and the Epanechnikov kernel. These have been succesfully used in the signal processing, density estimation, rendering, and inverse graphics literature (Liu et al., 2021; Jakob et al., 2011; Kerbl et al., 2023). Figure 3 shows their shape; we compare their performance in different applications in Section 8. Both of them have closed-form solutions for integration along segments as described in this section.
Figure 3. Visual comparison between the 1D (insets) and 2D Gaussian and Epanechnikov kernels. We plot ±$\pm{}3\sigma$ support to visualize the main difference between them: while the Gaussian kernel has infinite support, the Epanechnikov is finite, and decays sharply to 0. We further analyze differences between these kernels in Section 8.
@@ 5.1. Gaussian Kernel
The Gaussian kernel $G(\mathbf{x})$ is defined as
(13) $G(\mathbf{x})=\frac{1}{(2\pi)^{\frac{3}{2}}(|\Sigma|)^{\frac{1}{2}}}e^{-\frac{1}{2}(\mathbf{x}-\bm{\mu})^{T}\Sigma^{-1}(\mathbf{x}-\bm{\mu})},$
where $\Sigma$ is the $3\times 3$ covariance matrix and $\bm{\mu}$ its mean. As opposed to other convex kernels, Gaussians have unlimited support, which would potentially require evaluating all primitives in the scene for every ray. In practice, similar to Kerlb et al. (2023), we bound the kernel´s support to ±$\pm 3\sigma$, covering 99.73% of the Gaussian’s density.
@@
We compute the closed-form transmittance of a Gaussian primitive $\mathcal{P}_{i}$ from point $\mathbf{x}_{t_{0}}$ to point $\mathbf{x}_{t_{1}}$ along the ray $\mathrm{r}(t)=\mathbf{x}_{0}+\omega\,t$, all of them defined in the local space of the Gaussian, by integrating the optical depth $\tau_{i}(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}})$. Plugging in Equation (13) into Equation (11) we get
(14) $\tau_{i}(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}})$ $=\sigma_{i}\int_{t_{0}}^{t_{1}}G_{i}(\mathbf{x}_{t})\text{d}{t}$
$=\frac{\sigma_{i}\,e^{-\mathcal{C}_{1}}{2\mathcal{C}_{0}}\left(\operatorname{erf}\left(\frac{t_{0}(\mathcal{C}_{0})+\mathcal{C}_{2}}{(2\mathcal{C}_{0}|\Sigma|)^{\frac{1}{2}}}\right)-\operatorname{erf}\left(\frac{t_{1}(\mathcal{C}_{0})+\mathcal{C}_{2}}{(2\mathcal{C}_{0}|\Sigma|)^{\frac{1}{2}}}\right)\right)}{4\pi\sqrt{\mathcal{C}_{0}}},$
with $\operatorname{erf}(\cdot)$ being the error function, and where $\mathcal{C}_{0}$, $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are constants that depend on the ray position and direction, and the eigenvalues of the Gaussian’s covariance $\Sigma$. Explicit expressions for these constants can be found in the Appendix. Assuming exponential media, we can compute the transmittance as $T(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}})=\exp(-\tau_{i}(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}}))$.
@@
For sampling the free-flight distance $t$ in a single Gaussian primitive $\mathcal{P}_{i}$ with PDF $p(t)=\sigma_{i}G_{i}(\mathbf{x}_{t})\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})$, we need to invert its optical depth (14), which by setting $t_{0}=-\infty$ and using a random value $\xi\in(0,1)$ has closed form following
(15) $t(\xi)=\frac{\sqrt{2}\operatorname{erf}^{-1}(\mathcal{C}_{s}(\xi))(|\Sigma|^{\frac{1}{2}}\sqrt{\mathcal{C}_{0}})-\mathcal{C}_{2}}{\sigma_{i}\mathcal{C}_{0}},$
@@
For segments where more than one 3D Gaussian is contributing, we revert to using the bisection solver presented in Section 4. For single kernels, we normally use Equation 15. In practice, the piece-wise approximation of the $\operatorname{erf}^{-1}$ function commonly implemented in GPU math libraries (Kirk, 2007) produces large errors for certain input values, and for these we resort to the bisection solver as well, at the cost of some performance.
5.2. Epanechnikov Kernel
In our primitive-based rendering framework, Epanechnikov kernels have a significant advantage compared to Gaussian kernels due to their limited support. This allows more compact primitive shells, which accelerate rendering by reducing the number of overlaps. The 3D Epanechnikov kernel $\mathcal{E}(x)$ is defined as
(17) $\mathcal{E}(\mathbf{x})=\begin{cases}\frac{15}{8\pi(7^{3}\left|\Sigma\right|)^{\frac{1}{2}}}[1-\frac{1}{7}d(\mathbf{x})]&\text{ if }d(\mathbf{x})\leq 1\\ 0&\text{ otherwise },\end{cases}$
where, analogously to the Gaussian kernel, $\Sigma$ is the $3\times 3$ covariance matrix, $\bm{\mu}$ the mean, and $d(\mathbf{x})=(\mathbf{x}-\bm{\mu})^{T}\Sigma^{-1}(\mathbf{x}-\bm{\mu})$.
@@ 5.2. Epanechnikov Kernel
Analogously to the Gaussian kernel, we derive the closed-form transmittance of an Epanechnikov-based primitive between points $\mathbf{x}_{t0}$ and $\mathbf{x}_{t1}$ by integrating the optical depth between them as
@@
where $\mathcal{K}_{1}$, $\mathcal{K}_{2}$, $\mathcal{K}_{3}$ and $\mathcal{K}_{\text{norm}}$ are constants that depend on the ray position and direction, and the eigenvalues of the kernel’s covariance matrix $\Sigma$. Explicit expressions for these constants can be found in the Appendix. Analogously to any other kernel, and assuming exponential media, we can finally compute the transmittance as $T(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}})=\exp(-\tau_{i}(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}}))$.
@@
While there is an analytic solution for inverting Equation (12) for individual Epanechnikov kernels, it is not practical, given its complexity. Instead, we directly use the Newton-Raphson solver for segments with both one or multiple overlapping kernels.
@@ 6. Implementation Details
We implement our method in Mitsuba 3 (Jakob et al., 2022b), which we extended to support volumes of kernel primitives with different shell geometries, where we use ray tracing for querying primitives along the ray. We implement two different integrators, depending on the target application: 1) a volumetric-primitives path tracer (VPPT) supporting scattering media, and 2) a simplified volumetric-primitives radiance field (VPRF) integrator that computes the radiance field along a primary ray. In both cases, we compute transmittance and emission by following the segment-based formulation described in Section 4.2, by iterating over the segments in an ordered fashion: This fits very well with our ray-tracing-based querying of the primitives, since we collect new segments by simply casting new rays from the previous intersection. Listing 6.1 shows a skeleton pseudocode for our two integrators.
@@ 6.1. Integrators
This implements in our volumetric primitives framework the radiance field image formation model, similar to e.g. (Mildenhall et al., 2020; Kerbl et al., 2023). It significantly simplifies the image formation model from VPPT, by only considering the primary rays from the camera, accumulating radiance along the ray by querying a spherical harmonics-based emission in the primitives. In addition, we disable overlapping logic in order to favour speed. In practice, this means that we approximate Equation (7) using an ad-hoc approximate of the emission at each segment following
@@
{minted} [linenos,fontsize=, escapeinside=——, mathescape=true]python def primitive_tracing(ray, max_depth, P = []): depth = 0 —$t_{0}$— = 0.0 # Current segment start time while depth ¡ max_depth: V = [(p, p.—$t_{1}$—) for p in P] # List primitive exit points p = ray_intersect(ray) # Find next primitive V.append((p, p.—$t_{0}$—)) # Add primitive entry point # Process all exit points up to this intersection while not V.is_empty(): # Find next vertex to process v = min(V, key=lambda x: x.—$t$—) # Process segment according to application process_segment(ray, G, —$t_{0}$—, v.—$t$—) $t_{0}$— = v.—$t$— # Move on to next segment if p == v.p: break # Check if we have reached p P.remove(v.p) # Exiting v.p V.remove(v) P.append(p) # Ray is now entering primitive p ray = ray.move_to(p.—$t_{0}$—) # Update ray position depth += 1
{listing}
List of listings 0 Pseudo-code implementation of our primitive tracing algorithm. In VPPT, process_segment() samples a medium interaction in that segment and recursively calls primitive_tracing() on sampled position. In VPRF, process_segment() computes the radiance field contribution of that segment following Equation (21).

### v3

@@ 4. Volumetric Primitives
Inspired by recent works using Gaussian primitives for reconstructing radiance fields (kerbl20233Dgaussians), and following a similar medium definition as Knoll et al. (Knoll2021), we model media using sets of primitives. Each primitive $\mathcal{P}_{i}$ represents a statistical aggregate of matter with identical emission, cross section and phase function, and with density $\rho_{i}(\mathbf{x})$ defined by a three-dimensional un-normalized kernel $K_{i}(\mathbf{x})$. Given these primitives, we can model the distribution of matter, and therefore the extinction probability as a mixture of these volumetric primitives following
@@
with $N$ the number of primitives affecting $\mathbf{x}$, and $\sigma_{i}$ their cross-section. For kernels with infinite support, $N$ is the total number of primitives $N_{\text{total}}=N$. In our case, we assume that all kernels have limited support, so typically $N<N_{\text{total}}$. We compute the single-scattering albedo and phase function at $\mathbf{x}$ analogously, with the key difference of requiring normalization (see Appendix LABEL:app:opt_params).
@@ 4.1. The RTE with volumetric primitives
Defining matter as a mixture of volumetric primitives with finite support allow us to reinterpret Equation (3) as the summation of the contribution of all the primitives. For that, as shown in Figure 2 (bottom), we split the ray in segments based on the primitives’ entry and exit points; then, we integrate radiance along segments in a front-to-back pass, keeping track of the multiplicative transmittance which is computed analytically. This approach is similar to how production renderers deal with complex overlapping media (ProductionVolumeRendering2017, Ch.6); the key difference is that our approach approximates the media with primitives, which allows having closed-form transmittance and sampling expressions.
@@ 4.2. Solving Equation 7 for distance sampling
While Equation (7) is just a summation over segments that can be computed analytically (we show examples later in Section LABEL:sec:applications), in the most general case $L_{o,i}(\mathbf{x}_{t},\omega)$ hides the inscattering integral, which requires numerical evaluation using Monte Carlo sampling. Computing a Monte Carlo estimate would quickly become impractical; thus we need to sample a single segment per ray, which requires essentially sampling the distance $t$ along the ray, ideally with probability distribution function (PDF) $p(t)=\mu_{t}(\mathbf{x}_{t})\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})$.
While our primitive-based media allows the use of delta tracking (or any other Monte Carlo-based) estimators for distance sampling, we found that we can leverage the recursive formulation of transmittance (9) and pose the problem as an iterative search problem, where we uniformly sample the transmittance with a random variable $\xi\in(0,1)$, and search for the distance $t$ so that $\xi(t)=\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})$. This approach is very similar to regular tracking (stutton99regular), which we extend to handle overlapping kernels efficiently. In particular, we first search for a segment $\mathcal{S}_{k}$ so that $(1-\xi)\in[\textrm{T}_{k-1},\textrm{T}_{k}]$. Then, we invert the transmittance inside the segment, solving for $t\in[\widehat{t}_{k,0},\widehat{t}_{k,1}]$ the following equation:
@@
Depending on the type of kernel $K_{i}(\mathbf{x})$ being used, and the amount of overlaps in segment $\mathcal{S}_{k}$ this equation may have a closed-form analytical solution. However, for cases where Equation (12) does not have a simple analytical solution (e.g. many overlapping kernels at the same segment), we rely on numerical root-finding. In particular, we use two different methods depending on the kernel: the Newton-Raphson and the bisection solvers. The former is our primary choice, since it is efficient and converges with good precision with a few iterations (1-3), given that our problem is well-conditioned. However, for certain long-tailed kernels (e.g., Gaussian) it occasionally suffers from numerical instability. In these cases, we rely in the slightly slower but more stable bisection method. Details for both methods can be found in the Supplemental.
Biased uniform distance sampling inside segments
To maximize speed, in complex assets fitted with a high number of small kernels we can avoid inversion (12) and simply sample inside each segment uniformly. While this means we sample with pdf $p(t)\neq\mu_{t}(\mathbf{x}_{t})\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})$, in our implementation we assume that the probability distribution is close enough to it, allowing us to cancel out transmittance when computing the path throughput. This introduces a small amount of bias in exchange for a significant performance boost, which we analyze in Section LABEL:sec:forward_rendering.
@@ 5. Kernels
In the previous section, we have derived a general framework for radiative transport in primitive-based media, defined using arbitrary kernel functions. Any kernel could be used within our framework as long as 1) they have limited support or their decay is such that they can be bounded or clipped efficiently and 2) closed-form solutions to their line integrals exist or can be numerically computed efficiently.
Here we implement our framework using two different kernels: the Gaussian and the Epanechnikov kernel. These have been successfully used in the signal processing, density estimation, rendering, and inverse graphics literature (liu2021epan; jakob2011progressive; kerbl20233Dgaussians). Figure 3 shows their shape; we compare their performance in different applications in Section LABEL:sec:kernelcomp.
Figure 3. Visual comparison between the 1D (insets) and 2D Gaussian and Epanechnikov kernels. We plot $\pm{}3\sigma$ support to visualize the main difference between them: while the Gaussian kernel has infinite support, the Epanechnikov is finite, and decays sharply to 0. We further analyze differences between these kernels in Section LABEL:sec:kernelcomp.
@@ 5.1. Gaussian Kernel
The normalized Gaussian kernel $g_{i}(\mathbf{x})$ for primitive $\mathcal{P}_{i}$ is defined as
(13) $g_{i}(\mathbf{x})=\frac{1}{(2\pi)^{\frac{3}{2}}(|\Sigma_{i}|)^{\frac{1}{2}}}e^{-\frac{1}{2}(\mathbf{x}-\bm{c}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\bm{c}_{i})},$
where $\Sigma_{i}$ is the $3\times 3$ covariance matrix and $\bm{c}_{i}$ its mean. As opposed to other convex kernels, Gaussians have unlimited support, which would potentially require evaluating all primitives in the scene for every ray. In practice, similar to Kerlb et al. (kerbl20233Dgaussians), we bound the kernel’s support to three times the standard deviation, covering 99.73% of the Gaussian’s density.
@@
We compute the closed-form transmittance of a Gaussian primitive $\mathcal{P}_{i}$ from point $\mathbf{x}_{t_{0}}$ to point $\mathbf{x}_{t_{1}}$ along the ray $\mathrm{r}(t)=\mathbf{x}_{0}+\omega\,t$, by integrating the optical depth $\tau_{i}(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}})$. These parameters are them defined in the Gaussian’s local coordinate system, also known as whitened space (which we obtain by linearly transforming the ray into the canonical space of the Gaussian ellipsoidal shell, where the ellipsoid is a sphere of radius 1). Due to the symmetry of the Gaussian kernel around its mean, we can simplify the definite integral by shifting the origin of the ray to ensure a symmetric integration domain, integrating optical depth $\tau_{i}(\mathbf{x}_{-t_{s}},\mathbf{x}_{t_{s}})$ with $t_{s}=\frac{t_{0}+t_{1}}{2}$. This reduces the number of error functions required from 2 to 1. Plugging in Equation (13) into Equation (11) we get
(14) $\tau_{i}(\mathbf{x}_{-t_{s}},\mathbf{x}_{t_{s}})$ $=\sigma_{i}\int_{-t_{s}}^{t_{s}}g_{i}(\mathbf{x}_{t})\,\text{d}{t}=Ge^{-\frac{1}{2}\left(a-b^{2}\right)}\left[\operatorname{erf}\left(t_{s}\sqrt{\frac{1}{2}}\right)\right]$
with
(15) $a=\mathbf{x}_{0}\times\mathbf{x}_{0},\quad b=\mathbf{x}_{0}\times\omega,\quad G=\frac{\sigma_{i}}{2\pi\sqrt{v^{T}\Sigma_{i}^{-1}v|\Sigma_{i}|}}$
with $\operatorname{erf}(\cdot)$ being the error function. Assuming exponential media, we can compute the transmittance as $T(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}})=\exp(-\tau_{i}(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}}))$.
@@
For sampling the free-flight distance $t$ in a single Gaussian primitive $\mathcal{P}_{i}$ with PDF $p(t)=\alpha_{i}G_{i}(\mathbf{x}_{t})\textrm{T}(\mathbf{x}_{0},\mathbf{x}_{t})$, we need to invert its optical depth (14), which by setting $t_{0}=0$ and using a random value $\xi\in(0,1)$, and assuming an accumulated transmittance up to $t_{0}$ of $\beta$ has closed form following
(16) $t(\xi)=\sqrt{2}\operatorname{erf}^{-1}\left(\operatorname{erf}\left(\frac{b}{\sqrt{2}}\right)-\frac{2\operatorname{log}(\xi/\beta)}{Ge^{-\frac{1}{2}(a-b^{2})}}\right)-b$
For segments where more than one 3D Gaussian is contributing, we revert to using the bisection solver presented in Section 4. For single kernels, we normally use Equation 16. In practice, the piece-wise approximation of the $\operatorname{erf}^{-1}$ function commonly implemented in GPU math libraries (cuda) produces large errors for certain input values, and for these we resort to the bisection solver as well, at the cost of some performance.
5.2. Epanechnikov Kernel
In our primitive-based rendering framework, Epanechnikov kernels have a significant advantage compared to Gaussian kernels due to their limited support. This allows more compact primitive shells, which accelerate rendering by reducing the number of overlaps. The 3D Epanechnikov kernel $\mathcal{E}_{i}(x)$ for primitive $\mathcal{P}_{i}$ is defined as
@@ 5.2. Epanechnikov Kernel
where, analogously to the Gaussian kernel, $\Sigma_{i}$ is the $3\times 3$ covariance matrix, $\bm{c}_{i}$ the mean, and $d(\mathbf{x})=(\mathbf{x}-\bm{c}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\bm{c}_{i})$.
@@
We derive the closed-form transmittance of an Epanechnikov-based primitive between points $\mathbf{x}_{t_{0}}$ and $\mathbf{x}_{t_{1}}$ by integrating the optical depth between them as
@@
where $\mathcal{K}_{1}$, $\mathcal{K}_{2}$, $\mathcal{K}_{3}$ and $\mathcal{K}_{\text{norm}}$ are constants that depend on the ray position and direction, and the eigenvalues of the kernel’s covariance matrix $\Sigma_{i}$. Explicit expressions for these constants can be found in Appendix LABEL:app:kernel_transmittance. Analogously to any other kernel, and assuming exponential media, we can finally compute the transmittance as $T(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}})=\exp(-\tau_{i}(\mathbf{x}_{t_{0}},\mathbf{x}_{t_{1}}))$.
@@
While there is an analytic solution for inverting Equation (12) for one Epanechnikov kernel, it is not practical, given its complexity. Instead, we directly use the Newton-Raphson solver for segments with both one or multiple overlapping kernels.
@@ 6. Implementation Details
We implement our method in Mitsuba 3 (jakob2022mitsuba3), which we extended to support volumes of kernel primitives with different shell geometries, where we use ray tracing for querying primitives along the ray. We implement two different integrators, depending on the target application: 1) a volumetric-primitives path tracer (VPPT) supporting scattering media, and 2) a simplified volumetric-primitives radiance field (VPRF) integrator that computes the radiance field along a primary ray.
In both cases, we compute transmittance and emission by following the segment-based formulation described in Section 4.2, by iterating over the segments in an ordered fashion: This fits very well with our ray-tracing-based querying of the primitives, since we collect new segments by simply casting new rays from the previous intersection. Listing LABEL:code:primitive-tracing shows a skeleton pseudocode for our two integrators.
@@ 6.1. Integrators
This implements in our volumetric primitives framework the radiance field image formation model, similar to e.g. (mildenhall2020; kerbl20233Dgaussians). It significantly simplifies the image formation model from VPPT, by only considering the primary rays from the camera, accumulating radiance along the ray by querying a spherical harmonics-based emission in the primitives. In addition, we disable overlapping logic in order to favour speed. In practice, this means that we approximate Equation (7) using an ad-hoc approximate of the emission at each segment following